A 2.5 g sample of citric acid (below) is dissolved in water to give a 100 ml solution. Calculate the volume of 250 mM NaOH solution that would have to be added to neutralise 25 ml of the citric acid solution. 


1
Expert's answer
2021-01-18T01:58:00-0500

Molarity of citric acid is calculated using the following formula

"Molarity=\dfrac{moles}{volume}"


Number of moles of citric acid is given by;



"Moles=\dfrac{2.5}{192.124}"


"=0.013moles"


"Molarity=\dfrac{0.013}{100}=0.00013"